OpenStudy (anonymous):

Please help? Three times the least of three consecutive odd integers is three greater than two times the greatest. Find the greatest of the three integers.

OpenStudy (mathmale):

We need to represent these facts algebraically. I'd suggest that we let x represent the first odd integer; then the second odd integer will be represented by what? and the third odd integer will be represented by what? Hint: think about what "consecutive odd integer" means. What do consecutive odd integers look like when you write them out?

OpenStudy (anonymous):

Well, I understand that consecutive odd integers are like 3, 5, and 7. I know that this needs to be represented algebraically like you said, but I don't understand how to do that, even with the information you gave me.

OpenStudy (anonymous):

Would it be like 3x=3+2z? Where does the second integer fit in?

OpenStudy (mertsj):

x= first integer which is the smallest and is referred to as "the least" x+2=next odd integer x+4 = third odd integer 3x is "3 times the least" and that is 3 more than 2 times the biggest one So we have: 3x=2(x+4)+3

OpenStudy (anonymous):

Thank you! That makes perfect sense :)

OpenStudy (mertsj):

yw

OpenStudy (mathmale):

Cool. So, just solve 3x=2(x+4)+3 for x. Be certain to check you answer.

OpenStudy (anonymous):

Yeah, I got it. The greatest integer is 15 :)

OpenStudy (mertsj):

That is correct.

OpenStudy (mathmale):

Very glad for you.

OpenStudy (mertsj):

And notice that 3(11) is indeed 3 more than 2(15)

OpenStudy (anonymous):

:D